Grade 11Algebra

In a geometric progression, the first, third and fifth terms may be considered as the first, fourth, and sixteenth terms of an arithmetic progression. Determine the fourth term of this arithmetic progression knowing that its first term is 5.

Daksh Khetan
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
a = 5 given.
and 
ar^2 = a + 3d...............(1)
and
ar^4 = a + 15d.................(2)
solve eqns. 1 and 2 after substituting a=5
and write T4 = a+3d.